###############recursion (Python)
##def facto( num ):
## if num == 0:
## return 1
## else:
## return num * facto( num - 1 )
##
##Task 2. N! = 1*2*3*...*(N-1)*N as we know. There is another "double factorial":
##N!! :
##1)if N is odd then N!! = 1*3*5*7*9*...*N
##2)if N is even then N!! = 2*4*6*8*...*N
##
####Task 2. Create a function facto2 which calculates facto2(N) = N!!
##Show that your function works.
######use input(...)
Answer #1
#############recursion (Python)
def facto(num):
    if num == 0:
        return 1
    else:
        return num * facto(num - 1)


# Task 2. N! = 1*2*3*...*(N-1)*N as we know. There is another "double factorial":
# N!! :
# 1)if N is odd then N!! = 1*3*5*7*9*...*N
# 2)if N is even then N!! = 2*4*6*8*...*N

##Task 2. Create a function facto2 which calculates facto2(N) = N!!
####use input(...)
def facto2(n):
    if n <= 1:
        return 1
    else:
        return n * facto2(n - 2)


print(facto2(10))  # prints 10*8*6*4*2
print(facto2(9))  # prints 9*7*5*3*1
n = int(input("Enter a value for n: "))
print("facto2({}) = {}".format(n, facto2(n)))